OpenStudy (anonymous):

A manufacturer of skis offered chain discounts of 10/5/4 to many of its customers. Joe Jones ordered skis that had a total list price of $149,500. The amount of the trade discount is: a. $122,709.60 b. $26,970.40 c. $26,790.40 d. $122,907.69 e. None of these

OpenStudy (anonymous):

anser is c

OpenStudy (anonymous):

first take 149,500 * .10 = 14,950 149,500 - 14,950 = 134,550

OpenStudy (anonymous):

then, 134,550 * .05 = 6727.50 134,550 - 6727.50 = 127,822.50 127,822.50 * .04 = 5112.90

OpenStudy (anonymous):

trade discount = 14950 +6727.50+5112.90 = 26790.40